The energy for vacancy formation in BCC iron is 2.2 × 10^-19 J·mol^-1·K^-1. Calculate the fraction of lattice sites that would be vacant in 25 degree Celsius.
Answer #1

Ans;- Here Given,

temperature = 250C = 25+273=298K

Energy for vacancy formation , Ev = 2.2x10-19J/mol/K

Boltzmann constant , k = 1.38x10-23 J/molK

Let fv be the fraction of lattice sites ,

therefore

f_{v}=exp(\frac{-E_{v}}{kT})

      = exp[(-2.2x10-19J/mol/K ) / 1.38x10-23 J/molK x298K]

      =exp(-53.496)

      = 5.847x10-24